OpenStudy (anonymous):

What do the following two equations represent? 5x+4y=4 −8x+10y=5 perpendicular, parallel equivalent none of the above

OpenStudy (anonymous):

pretty sure it's perpindicular

OpenStudy (anonymous):

youre right!

OpenStudy (anonymous):

mind helping with another

OpenStudy (anonymous):

can you give me best response

OpenStudy (anonymous):

and yeah sure

OpenStudy (anonymous):

Whats the question?

OpenStudy (anonymous):

Are you there?

OpenStudy (anonymous):

What do the following two equations represent? 4x+2y=2 −12x−6y=−6